A solenoid of length 2.2m has a diameter of 0.9. The magnetic field at its centre is 0.4tesla . estimate the energy stored in the field of this coil.

1
Expert's answer
2022-02-28T13:51:29-0500

Gives

"B=0.4T"

"l=2.2m"

"r=\frac{0.9}{2}" m

"U=\frac{B^2}{2\mu}Al"

Area

"A=\pi r^2"


"U=\frac{0.4^2}{2\times4\pi\times10^{-7}}\times\frac{\pi \times0.81\times2.2}{4}=89100J"